Green Steering Mirror Update
Yesterday, Nick and I completed the green steering mirrors upgrade. I attached the file that contained the procedure that we plan before we did the upgrade. We placed an iris at the input of the OL and we place another iris before the harmonic separator. We did not use the beam scanner because someone was using it, so what we did was to assume that the cavity is well align and place the iris so that we can recover the alignment. We used the measuring tape to approximate as close as we could the position where the lenses were supposed to go. I did a measurement of the derivative of the waist size in terms of the position of the lens and the derivative of the waist Position in terms of the lenses position at the optimum solution that a la mode give us. Because of this plot, we decide to mount lens 3 and lens 5 into translational stages. After mounting each lenses and mirrors we worked on the alignment of the beam into the cavity. We were able to align the green into the cavity and we were able to locked the cavity to the TEM00 mode. We started to work on the optimization of the mode matching. However, the maximum mode matching that we got was around 0.6, which we need to work a little bit more on the tuning of the mode matching. We leave the iris mounted on the table. I took a picture of the table, and I attached below. For the OL, we just make sure that the output where somehow hitting the QPD, but we didn't really I aligned it. We need to work a little bit more on the alignment of the OL and the tuning of the mirror to maximize the green mode matching. |